# PO-Flow: Flow-based Generative Models for Sampling Potential Outcomes and Counterfactuals

### v2

@@ 4.  Set intervened action to 1a1-a;
Counterfactual Prediction
Algorithm 2
Counterfactual Prediction
Because of the smooth and information-preserving nature of the continuous normalizing flow path, the encoded $z$ retains rich information from the factual outcome $y^{(a)}$, enabling accurate decoding of the counterfactual $\hat{y}_{\text{cf}}^{(1-a)}$ under counterfactual treatment 1a1-a. To our knowledge, this work is among the early attempts at counterfactual prediction in the Potential Outcomes (POs) literature.
In the following, we present a supporting result on the counterfactual recovery properties of PO-Flow under monotone structural causal models (SCMs). While PO-Flow does not explicitly aim to model the true SCM during training or inference, we show that if the data are generated from monotone SCMs, then PO-Flow provides recovery guarantees under certain conditions.
We denote the SCM for an individual as $Y^{(A)}=f_{X}(A,U)$, where $U$ denotes the exogenous noise. The encoding function is given by $Z:=\Phi_{\theta}(Y^{(A)};\,X,A)$, which corresponds to Equation (10). We begin by introducing the following assumptions:
Assumption 4.1.
(A1) $U\perp\!\!\!\perp A$.
(A2) The structural causal model $f_{X}(A,U)$ is monotone in $U$ for each action $A\in\{0,1\}$.
(A3) For every $x$, the encoded latent $Z=\Phi_{\theta}(Y^{(A)};\,X,A)$ induced by the flow is statistically independent of $A$, i.e., $Z\perp\!\!\!\perp A\ |\ X=x$.
Remark 4.2.
@@
In the POs framework, covariates $x$ are multivariate while the potential outcome is univariate, $y\in\mathbb{R}$. Therefore, (A2) is invoked for the univariate case. It is automatically satisfied under an additive structural causal model, $Y^{(A)}=f_{X}(A,U):=f_{X}^{*}(A)+U$. Under certain identifiability conditions, it also holds for non-linear models [62, 55]. The following theoretical results hold only under monotone SCM settings.
@@
(A3) assumes the encoded representation $Z=\Phi_{\theta}(Y^{(A)};\,X,A)$ in PO-Flow is statistically independent of treatment $A$, though not necessarily functionally invariant to $A$. In the infinite-data limit with exact training, the continuous normalizing flow (Equation (8)) maps each treatment-conditioned outcome $Y$ to the same base distribution $q(Z)=\mathcal{N}(0,1)$, so $p_{\theta}(Z\mid X,A)=q(Z)$ and (A3) holds exactly. In finite samples, deviations may arise; we assess this empirically in Appendix D.1.
Under the Assumption 4.1, we present a result for PO-Flow when the true SCM is monotone:
Proposition 4.3 (Encoded $Z$ as a function of the exogenous noise $U$).
@@
Let Assumption 4.1 hold. Assume that the exogenous noise $U\sim\mathrm{Unif}[0,1]$. Then there exists a continuously differentiable bijection function $\psi_{X}:\;\mathcal{U}\;\longrightarrow\;\mathcal{Z}$ that is independent of the treatment assignment $A$, i.e.,
$\small Z\;=\;\Phi_{\theta}(Y^{(A)};X,A)\;=\;\Phi_{\theta}(f_{X}(A,U);X,A)\;=\;\psi_{X}\bigl(U\bigr)$ (12)
Remark 4.4.
We can transform the exogenous noise $U\sim\mathrm{Unif}[0,1]$ assumed in Proposition 12 into other distributions, such as Gaussian. If $Z\sim\mathcal{N}(0,1)$ with CDF $F(Z)$, then $U=F(Z)\sim\mathrm{Unif}[0,1]$, and any structural assignment $f(\cdot,U)$ can be written equivalently as $\tilde{f}(\cdot,Z)=f(\cdot,F(Z))$. Therefore, the conclusions of Proposition 12 also hold with other noise distributions.
@@
Following Proposition 12, we have the following result on counterfactual recovery under monotone SCMs:
Corollary 4.5 (Counterfactual recovery under monotone SCMs).
@@
Assume the conditions in Assumption 4.1 hold. Consider a factual sample $\bigl(Y^{(A)},X,A\bigr)$ given by the structural equation $Y^{(A)}=f_{X}\bigl(A,U\bigr)$, and let its encoded representation be $Z=\Phi_{\theta}(Y^{(A)};X,A)$. Apply the intervention $\mathrm{do}\bigl(A=\gamma\bigr)$. Assume that the true counterfactual is given by $Y_{\mathrm{CF}}^{(\gamma)}=f_{X}(\gamma,U)$. The decoder then recovers the counterfactual almost surely:
Remark 4.6.
One may notice that Assumptions (A1) and (A2) mirror those in Bijective Generation Mechanisms (BGM) [43]. BGM establishes class-level identifiability: if a learned mechanism lies in the BGM class and matches the observational distribution ($\hat{f}_{X}(A,U)\stackrel{{\scriptstyle d}}{{=}}f_{X}(A,U)$), then counterfactuals are identifiable. By contrast, our Prop. 4.3 shows that PO-Flow’s encoder is bijective in $U$, placing our model within the BGM class. We then use an alternative constructive route (building on the intermediate step in Prop. 4.3) to obtain Cor. 4.5 under (A1)–(A3). Alternatively, one could invoke the BGM result after Prop. 4.3 to reach Cor. 4.5; see Appendix C.
4.3 PO-Flow for Learning Potential Outcomes Densities
Another advantage of PO-Flow is its ability to learn the density of the Potential Outcomes (POs). We define $p_{\theta,Y|X,A}(\cdot\mid x,a,z)$ as the PO density computed by PO-Flow, conditioned on $(x,a)$ and a base sample $z$. The explicit form of this density, mapping from latent to data space, is presented in Proposition 4.7, with the proof provided in Appendix A.
Proposition 4.7.
Given a noise sample $z\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
Empirically, the divergence of the velocity field can be computed much more efficiently under the Hutchinson trace estimator (e.g., [27, 58]). More details are provided in Appendix B.4.
Unlike prior works [40, 41] that generate PO samples and approximate empirical distributions, PO-Flow provides direct likelihood evaluation. This allows principled uncertainty quantification, supports ranking outcomes by plausibility (Table 4), enables information-theoretic analyses such as KL divergence (Table 3), and facilitates tasks like anomaly detection via quantile-based thresholds.
5 Experiments
Baselines: Our baselines cover several categories: model-agnostic learners that learn the conditional distribution means (S-learner [33], T-learner [33], and DR-Learner [32]), representation-based conditional regressors (Counterfactual Regression (CFR) [50]), deep generative methods (Causal Effect Variational Autoencoder (CEVAE) [38]; GAN for Individualized Treatment Effect (GANITE) [61]), and most recent generative approaches based on diffusion (DiffPO [39]) and discrete normalizing flows (INFs [41]).
Datasets: Counterfactual outcomes and unobserved potential outcomes (POs) are inaccessible in real-world data but can be simulated in semi-synthetic settings, as discussed in detail in Appendix B.1. Accordingly, we evaluate PO-Flow on synthetic benchmarks, including ACIC 2016, ACIC 2018, IHDP, and IBM.
The code and datasets are included in the supplementary materials for anonymous submission.
5.1 Prediction of Potential Outcomes
Table 1: Root mean squared error (RMSE) of estimated potential outcomes (POs) across different methods on the ACIC 2018, IHDP, and IBM datasets. Results are averaged over 10-fold cross-validation.
ACIC 2018 IHDP IBM
PO-Flow $\mathbf{0.53_{\pm.10}}$ $0.60_{\pm.12}$ $\mathbf{0.98_{\pm.11}}$ $\mathbf{1.19_{\pm.13}}$ $\mathbf{0.90_{\pm.09}}$ $\mathbf{0.97_{\pm.11}}$
INFs $0.69_{\pm.16}$ $0.78_{\pm.17}$ $1.02_{\pm.10}$ $\mathbf{1.20_{\pm.12}}$ $1.52_{\pm.24}$ $1.57_{\pm.28}$
S-learner $\mathbf{0.54_{\pm.09}}$ $\mathbf{0.57_{\pm.11}}$ $1.31_{\pm.18}$ $1.44_{\pm.20}$ $01_{\pm.17} 1.1$$16_{\pm.19}$
CEVAE $0.83_{\pm.17}$ $0.85_{\pm.17}$ $1.18_{\pm.15}$ $1.36_{\pm.18}$ $0.96_{\pm.12}$ $\mathbf{0.98_{\pm.12}}$
@@ 5.1 Prediction of Potential Outcomes
Evaluation Metrics: We evaluate the point estimates of potential outcomes (POs) using the root mean squared error (RMSE) between the estimated outcomes $\hat{y}_{i}^{(a)}$ and the true outcomes $y_{i}^{(a)}$, computed as $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}^{(a)}-\hat{y}_{i}^{(a)})^{2}}$, where $\text{RMSE}_{\text{in}}$ and $\text{RMSE}_{\text{out}}$ refer to the RMSE evaluated on the training and test sets, respectively.
@@
Results: Table 1 shows that PO-Flow achieves the most accurate prediction of potential outcomes on the majority of datasets.
Besides, we also report in Table 4 the RMSE of the most likely POs selected by PO-Flow using the highest log-density (as computed in Proposition 4.7).
5.2 Estimation of CATEs
Conditional Average Treatment Effect (CATE): The CATE is defined as $\tau(x)=\mathbb{E}[y^{(a=1)}-y^{(a=0)}\mid X=x]$, representing the expected treatment effect conditioned on covariates $x$. It captures individual-level heterogeneity in causal effects.
Evaluation Metrics: We evaluate CATE using the precision in estimation of heterogeneous effect (PEHE), defined as the mean squared error between estimated and true effects: $\epsilon_{\text{PEHE}}=\frac{1}{n}\sum_{i=1}^{n}(\hat{\tau}(x_{i})-\tau(x_{i}))^{2}$. Table 2 reports the root PEHE for both training and test sets. PO-Flow achieves the most accurate CATE estimates on most datasets.
Table 2: Root Precision in Estimation of Heterogeneous Effect ($\sqrt{\epsilon_{\text{PEHE}}}$) across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets.
ACIC 2016 ACIC 2018 IHDP IBM
$\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{in}}}$ $\sqrt{\epsilon_{\text{PEHE}}^{\text{out}}}$
PO-Flow $\mathbf{0.76_{\pm.08}}$ $\mathbf{0.82_{\pm.09}}$ $\mathbf{0.05_{\pm.01}}$ $\mathbf{0.07_{\pm.02}}$ $\mathbf{0.41_{\pm.08}}$ $\mathbf{0.45_{\pm.10}}$ $\mathbf{0.04_{\pm.01}}$ $\mathbf{0.04_{\pm.01}}$
[…]
INFs $1.37_{\pm.28}$ $1.42_{\pm.30}$ $0.33_{\pm.11}$ $0.37_{\pm.12}$ $0.90_{\pm.27}$ $0.95_{\pm.27}$ $0.16_{\pm.08}$ $0.18_{\pm.08}$
S-learner $3.81_{\pm.76}$ $3.87_{\pm.98}$ $\mathbf{0.06_{\pm.02}}$ $\mathbf{0.07_{\pm.03}}$ $0.92_{\pm.07}$ $93_{\pm.08} 0.0$$6_{\pm.02} 0.0$$07_{\pm.04}$
T-learner $5.02_{\pm.71}$ $5.15_{\pm.87}$ $1.21_{\pm.47}$ $1.32_{\pm.53}$ $1.24_{\pm.08}$ $1.31_{\pm.09}$ $0.60_{\pm.14}$ $0.69_{\pm.15}$
DR-learner $3.04_{\pm.19}$ $3.51_{\pm.23}$ $0.43_{\pm.08}$ $0.57_{\pm.10}$ $1.03_{\pm.09}$ $1.28_{\pm.15}$ $0.12_{\pm.03}$ $0.19_{\pm.05}$
CEVAE $2.06_{\pm.61}$ $2.31_{\pm.74}$ $1.19_{\pm.15}$ $1.21_{\pm.16}$ $1.41_{\pm.10}$ $1.43_{\pm.10}$ $0.73_{\pm.15}$ $0.75_{\pm.16}$
CFR $2.80_{\pm.58}$ $2.74_{\pm.60}$ $0.77_{\pm.18}$ $0.80_{\pm.18}$ $0.73_{\pm.04}$ $0.78_{\pm.04}$ $1.05_{\pm.27}$ $1.08_{\pm.28}$
GANITE $3.02_{\pm.76}$ $3.11_{\pm.74}$ $0.55_{\pm.11}$ $0.60_{\pm.13}$ $1.90_{\pm.40}$ $2.40_{\pm.40}$ $0.77_{\pm.20}$ $0.80_{\pm.24}$
5.3 Learning the Full Potential Outcome Distributions
Beyond point estimation, PO-Flow offers two distinct advantages for modeling potential outcomes: (1) it enables individualized sampling from the conditional distribution $P(Y\mid X,A)$, generating a rich set of plausible outcomes for each individual; and (2) it supports learning the density of any generated outcome, providing deeper insights into predictive uncertainty.
5.3.1 Learning potential outcome densities
PO-Flow enables direct computation of potential outcome densities, as shown in Proposition 4.7. To evaluate the learned densities, we compute the KL divergence between the PO-Flow estimate $p_{\theta,Y|X,A}(\cdot|x,a)$ and the ground-truth density $p_{\text{true}}(\cdot|x,a)$. Given that the semi-synthetic datasets assume $y^{(a)}\sim\mathcal{N}(\mu(a),1)$, the individual-level KL is $\text{KL}(\hat{Y}_{X}^{(a)}\|Y_{X}^{(a)})=\mathbb{E}_{\hat{y}\sim p_{\theta,Y|X,A}}\left[\log\frac{p_{\theta,Y|X,A}(\hat{y}|x,a)}{\mathcal{N}(\hat{y};\mu(a),1)}\right]$, where $p_{\theta,Y\mid X,A}$ is computed via (14).

### v3

@@ 4.  Set intervened treatment to 1a1-a;
Algorithm 2 Counterfactual Prediction
IV Proposed Method
This section presents PO-Flow, a continuous normalizing flow framework that jointly supports (i) individualized potential outcome prediction, (ii) factual-conditioned counterfactual prediction, and (iii) learning potential outcome densities.
We model outcomes using a continuous normalizing flow whose velocity field $v_{\theta}$ is conditioned on covariates $x$ and treatment $a$:
$\frac{dy(t)}{dt}=v_{\theta}\!\bigl(y(t),t;\,x,a\bigr),\qquad t\in[0,1].$ (5)
Given a data point $(y^{(a)},x,a)\sim p_{Y,X,A}$, we set $y_{0}=y^{(a)}$ as the observed training sample at time $t=0$ and draw a base sample $y_{1}=z\sim q(\cdot)$ at $t=1$.
Under this conditional formulation, the training objective retains the form of conditional flow matching (4) and is given by
$\mathcal{L}_{\mathrm{CFM}}(\theta)=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ (y^{(a)},x,a)\sim p_{Y,X,A},\\ y_{1}\sim q(\cdot)\end{subarray}}\left\|v_{\theta}(\phi_{t},t;\,x,a)-\frac{d\phi_{t}}{dt}\right\|^{2},$ (6)
where $\phi_{t}$ and $d\phi_{t}/dt$ are defined in (2) and (3), respectively.
Remark IV.1.
Following prior work [22], one may estimate propensity scores $w(x)=p(A=1\mid X=x)$ and apply inverse weighting $1/w(x)$ in the loss. We observed limited empirical improvement and therefore include both weighted and unweighted variants in our implementation.
IV-A Encoding and Decoding via CNFs
We interpret the forward process as an encoding operation, denoted by $Z:=\Phi_{\theta}(Y^{(A)};\,X,A)$:
$z:=\Phi_{\theta}(y^{(a)};\,x,a)=y^{(a)}+\int_{0}^{1}v_{\theta}(y(t),t;\,x,a)\,dt,$ (7)
with $y(0)=y^{(a)}$.
Conversely, the reverse process serves as a decoding operation:
$\hat{y}_{0}:=\Phi_{\theta}^{-1}(z;\,x,a)=z-\int_{0}^{1}v_{\theta}(y(t),t;\,x,a)\,dt,$ (8)
with $y(1)=z$.
To generate a potential outcome $\hat{y}^{(a)}$ for an individual with $(x,a)$, we sample $z\sim q(\cdot)$ and decode it using (8). Algorithm 1 summarizes the procedure.
Counterfactual prediction follows an encode–decode procedure. Given a factual sample $(y^{(a)},x,a)$, the forward process (7) encodes it into $z$ which captures the factual information, and the reverse process (8) then decodes $z$ under the counterfactual treatment 1a1-a to obtain $\hat{y}_{\mathrm{cf}}^{(1-a)}$. Algorithm 2 summarizes the procedure.
V Theoretical Properties
We characterize theoretical properties of PO-Flow for (i) factual-conditioned counterfactual recovery under monotone SCMs and (ii) likelihood evaluation for potential outcomes.
V-A Counterfactual Recovery under Monotone SCMs
We assume the underlying data-generating process follows a structural causal model (SCM)
$Y^{(A)}=f_{X}(A,U),$
where $U$ denotes exogenous noise. PO-Flow defines an encoding map $Z:=\Phi_{\theta}(Y^{(A)};X,A)$ via (7).
Assumption V.1.
  1. (A1)
    $U\perp\!\!\!\perp A$.
  2. (A2)
    For each $A\in\{0,1\}$, the SCM $f_{X}(A,U)$ is strictly monotone in $U$.
  3. (A3)
    The encoded latent is conditionally treatment-invariant: $Z\perp\!\!\!\perp A\mid X$.
Remark V.2.
@@ V-A Counterfactual Recovery under Monotone SCMs
In the potential outcomes setting, $X$ is multivariate while the outcome is univariate ($y\in\mathbb{R}$), so (A2) concerns monotonicity in the scalar noise $U$. It holds automatically under additive SCMs $Y^{(A)}=f_{X}^{*}(A)+U$, and can also hold under certain identifiability conditions for nonlinear models [39, 34].
Remark V.3.
@@
Assumption (A3) posits that the encoding $Z=\Phi_{\theta}(Y^{(A)};X,A)$ is conditionally independent of treatment given $X$. In the infinite-data limit with exact training, the CNF maps each treatment-conditioned outcome distribution to the same base distribution $q(Z)$, yielding $p_{\theta}(Z\mid X,A)=q(Z)$ and (A3) is expected to hold. In finite samples, deviations may arise; we assess this empirically in Appendix D.1 of [35].
Under Assumption V.1, we obtain the following result; see Appendix A of [35] for proofs.
Proposition V.4 (Encoded $Z$ as a function of exogenous noise $U$).
@@
Let Assumptions III.1 and V.1 hold and assume $U\sim\mathrm{Unif}[0,1]$. Then there exists a continuously differentiable bijection $\psi_{X}$ independent of AA such that
$Z=\Phi_{\theta}(Y^{(A)};X,A)=\Phi_{\theta}(f_{X}(A,U);X,A)=\psi_{X}(U).$
Remark V.5.
The assumption $U\sim\mathrm{Unif}[0,1]$ is without loss of generality: if $Z\sim\mathcal{N}(0,1)$ with CDF $F$, then $U=F(Z)\sim\mathrm{Unif}[0,1]$, and the SCM can be equivalently written in terms of $Z$.
@@
Following Proposition V.4, we have the following result on counterfactual recovery under monotone SCMs; see Appendix A of [35] for proofs.
Corollary V.6 (Counterfactual recovery under monotone SCMs).
@@
Assume Assumptions III.1 and V.1 hold. Consider a factual sample $(Y^{(A)},X,A)$ generated by $Y^{(A)}=f_{X}(A,U)$ and its encoding $Z=\Phi_{\theta}(Y^{(A)};X,A)$. Under the intervention $\mathrm{do}(A=\gamma)$, the true counterfactual is $Y_{\mathrm{cf}}^{(\gamma)}=f_{X}(\gamma,U)$, and the PO-Flow decoder recovers it almost surely:
$\hat{Y}_{\mathrm{cf}}^{(\gamma)}:=\Phi_{\theta}^{-1}(Z;X,\gamma)=Y_{\mathrm{cf}}^{(\gamma)}.$
V-B Likelihood Evaluation for Potential Outcome Densities
PO-Flow also enables explicit likelihood evaluation for potential outcomes via the change-of-variables formula for CNFs; see Appendix A of [35] for the proof.
Proposition V.7.
Given a base sample $z\sim q(\cdot)$ at $t=1$, the log-density of the decoded outcome at $t=0$ is
@@ V-B Likelihood Evaluation for Potential Outcome Densities
where $y=\Phi_{\theta}^{-1}(z;x,a)$.
VI Experiments
We compare PO-Flow with model-agnostic mean learners (S-, T-, and DR-learners [18, 17]), representation-based regression (CFR [30]), deep generative methods (CEVAE [21], GANITE [38]), and recent generative baselines based on diffusion (DiffPO [22]) and discrete normalizing flows (INFs [24]). For likelihood-based comparisons, we additionally include Gaussian-process baselines for interventional treatment effects (GP-ITE) [3, 13]. Since counterfactuals and unobserved potential outcomes are unavailable in real-world data, we evaluate on semi-synthetic benchmarks: ACIC 2016/2018, IHDP, and IBM. All results are averaged over 10-fold cross-validation; in all tables, columns I and O denote in-sample (training) and out-of-sample (test) performance, respectively.
We evaluate PO-Flow across five complementary tasks, each designed to assess a distinct capability of the proposed framework. These include (i) point prediction of potential outcomes, (ii) conditional average treatment effect estimation, (iii) learning potential outcome densities for likelihood-based evaluation, (iv) distributional fit of potential outcomes, and (v) factual-conditioned counterfactual outcome prediction. Together, these tasks reflect the core claims of PO-Flow and provide a comprehensive evaluation of its individual-level predictive, distributional, and counterfactual reasoning capabilities.
VI-A Task 1: Potential Outcome Prediction
@@
We evaluate point prediction of potential outcomes using RMSE using the standard definition $\mathrm{RMSE}=[\frac{1}{n}\sum_{i=1}^{n}(y_{i}^{(a)}-\hat{y}_{i}^{(a)})^{2}]^{1/2}.$ Table I shows that PO-Flow achieves the strongest performance on the majority of datasets.
TABLE I: RMSE of estimated potential outcomes (10-fold cross-validation). Parentheses report standard deviation. I/O denote in-/out-of-sample.
ACIC18 IHDP IBM
Method I O I O I O
PO-Flow 0.53(0.10) 0.60(0.12) 0.98(0.11) 1.19(0.13) 0.90(0.09) 0.97(0.11)
DiffPO 0.52(0.13) 0.59(0.15) 1.33(0.20) 1.40(0.22) 1.82(0.38) 1.85(0.35)
INFs 0.69(0.16) 0.78(0.17) 1.02(0.10) 1.20(0.12) 1.52(0.24) 1.57(0.28)
S-learner 0.54(0.09) 0.57(0.11) 1.31(0.18) 1.44(0.20) 1.01(0.17) 1.16(0.19)
T-learner 1.57(0.32) 1.71(0.40) 1.45(0.25) 1.49(0.27) 1.89(0.47) 1.96(0.50)
CEVAE 0.83(0.17) 0.85(0.17) 1.18(0.15) 1.36(0.18) 0.96(0.12) 0.98(0.12)
CFR 0.94(0.10) 0.98(0.12) 1.10(0.20) 1.18(0.20) 2.09(0.45) 2.17(0.48)
GANITE 0.88(0.15) 0.97(0.15) 1.60(0.36) 1.67(0.34) 2.48(0.39) 2.59(0.47)
VI-B Task 2: Conditional Average Treatment Effect (CATE)
CATE is defined as $\tau(x)=\mathbb{E}[y^{(1)}-y^{(0)}\mid X=x]$, representing the expected treatment effect conditioned on covariates $x$. We evaluate CATE using the precision in estimation of heterogeneous effect (PEHE) [1], defined as:
$\epsilon_{\text{PEHE}}=\frac{1}{n}\sum_{i=1}^{n}(\hat{\tau}(x_{i})-\tau(x_{i}))^{2},$ (10)
where $\hat{\tau}(x)=\mathbb{E}_{p_{\theta}(\cdot\mid x,1)}[Y]-\mathbb{E}_{p_{\theta}(\cdot\mid x,0)}[Y]$, approximated by Monte Carlo samples from PO-Flow. Table II reports the root PEHE for training and test sets. PO-Flow achieves the most accurate CATE estimates on most datasets.
TABLE II: Measuring precision in estimation of the heterogeneous effect using $\sqrt{\epsilon_{\text{PEHE}}}$; Standard deviations omitted and they are reasonably small. I/O denote in-/out-of-sample.
ACIC16 ACIC18 IHDP IBM
Method I O I O I O I O
PO-Flow 0.76 0.82 0.05 0.07 0.41 0.45 0.04 0.04
DiffPO 0.91 1.12 0.07 0.08 0.79 0.84 0.08 0.09
INFs 1.37 1.42 0.33 0.37 0.90 0.95 0.16 0.18
S-learner 3.81 3.87 0.06 0.07 0.92 0.93 0.06 0.07
T-learner 5.02 5.15 1.21 1.32 1.24 1.31 0.60 0.69
DR-learner 3.04 3.51 0.43 0.57 1.03 1.28 0.12 0.19
CEVAE 2.06 2.31 1.19 1.21 1.41 1.43 0.73 0.75
CFR 2.80 2.74 0.77 0.80 0.73 0.78 1.05 1.08
GANITE 3.02 3.11 0.55 0.60 1.90 2.40 0.77 0.80
VI-C Task 3: Learning Potential Outcome Densities
PO-Flow supports explicit likelihood evaluation of potential outcomes via Proposition V.7. We evaluate density accuracy using the KL divergence between the learned density $p_{\theta,Y\mid X,A}(\cdot\mid x,a)$ and the ground-truth density $p_{\mathrm{true}}(\cdot\mid x,a)$. For the semi-synthetic benchmarks, $Y^{(a)}\sim\mathcal{N}(\mu_{X}(a),1)$, and we compute
$\mathrm{KL}\!\left(\hat{Y}^{(a)}\,\|\,Y^{(a)}\right)=\mathbb{E}_{\hat{y}\sim p_{\theta,Y\mid X,A}}\left[\log\frac{p_{\theta,Y\mid X,A}(\hat{y}\mid x,a)}{\mathcal{N}(\hat{y};\mu_{X}(a),1)}\right],$
where $p_{\theta,Y\mid X,A}$ is evaluated via (9). As shown in Table III, PO-Flow achieves lower KL divergence than GP-ITE and INF on ACIC 2018, IHDP, and IBM.